OpenStudy (anonymous):

PLEASE give me guidance? What is the value of y in the solution to the following system of equations?5x - 3y = -32x - 6y = -6 Answers: -1 0 1 2

OpenStudy (goformit100):

-1

OpenStudy (anonymous):

How did you get that?

OpenStudy (anonymous):

And thank you for replying...

OpenStudy (anonymous):

Please help

OpenStudy (anonymous):

5x-3y=-6 => x=(3y-6)/5 -32x-6y=-6 => x=(6y-6)/-32 (3y-6)/5 =(6y-6)/-32 -96y+192 = 30y-30 126y=222 y=222/126 y=37/21 no answer

OpenStudy (anonymous):

There has to be one

OpenStudy (ash2326):

@InsanelyChaotic Are these your equations? \[5x-3y=-32\] \[x-6y=-6\]

OpenStudy (anonymous):

Check out..i too didn't get any answer from your choices